OpenStudy (anonymous):

f(x) = x^4 - x^2 - 30

OpenStudy (solomonzelman):

let x^2=a 0=a^2-a-30 (a-.5)^2=30 + .25 a-.5= +-5.5 a=.5+-5.5 a= -5 , 6

OpenStudy (solomonzelman):

so x^2=6 x= +-sqrt6 x^2=-5 x= +-isqrt5

OpenStudy (solomonzelman):

\[\sqrt{6},~~~~-\sqrt{6},~~~~i \sqrt {5},~~~~-i \sqrt {5}\]

OpenStudy (anonymous):

Wouldn't that just become x^2-30?

OpenStudy (solomonzelman):

no, it would be a^2-a-30

OpenStudy (anonymous):

Where are you getting that a from? o.O